### v2

# Generalized Logit Adjustment: Calibrating Fine-tuned Models by Removing Label Bias in Foundation Models

Abstract
Foundation models like CLIP allow zero-shot transfer on various tasks without additional training data. Yet, the zero-shot performance is less competitive than a fully supervised one. Thus, to enhance the performance, fine-tuning and ensembling are also commonly adopted to better fit the downstream tasks. However, we argue that such prior work has overlooked the inherent biases in foundation models. Due to the highly imbalanced Web-scale training set, these foundation models are inevitably skewed toward frequent semantics, and thus the subsequent fine-tuning or ensembling is still biased. In this study, we systematically examine the biases in foundation models and demonstrate the efficacy of our proposed Generalized Logit Adjustment (GLA) method. Note that bias estimation in foundation models is challenging, as most pre-train data cannot be explicitly accessed like in traditional long-tailed classification tasks. To this end, GLA has an optimization-based bias estimation approach for debiasing foundation models. As our work resolves a fundamental flaw in the pre-training, the proposed GLA demonstrates significant improvements across a diverse range of tasks: it achieves 1.5 pp accuracy gains on ImageNet, a large average improvement (1.4-4.6 pp) on 11 few-shot datasets, 2.4 pp gains on long-tailed classification. Codes are in https://github.com/BeierZhu/GLA.
1 Introduction
Thanks to the Web-scale data and self-supervised strategies, foundation models like CLIP radford2021learning empower zero-shot transfer to a wide variety of domains brown2020language ; alayrac2022flamingo ; yu2022coca . However, the zero-shot performance is still weak on several domain-specific tasks such as differentiating models of cars, species of flowers, and variants of aircraft radford2021learning ; brown2020language . Therefore, it is a common practice to improve the downstream performance via supervised fine-tuning on labeled data, e.g., linear probing, prompt tuning COOP ; prograd , and end-to-end fine-tuning.
However, fine-tuned models are easily biased: they are adept in exploiting spurious correlations that only hold on the downstream distribution radford2021learning ; pham2023combined ; wortsman2022robust ; prograd . To improve the robustness, several studies wortsman2022robust ; prograd ; zhu2023debiased propose to combine fine-tuned models with zero-shot models. For example, WiSE-FT wortsman2022robust ensembles the fine-tuned and zero-shot models in weight space and ProGrad prograd uses zero-shot predictions to regularize the fine-tuning gradient. The underlying assumption lies in that the zero-shot models are robust to distribution shifts radford2021learning , and their predictions are complementary to those of fine-tuned models wortsman2022robust .
Despite these methods exhibiting performance gains on both in-distribution and out-of-distribution evaluations, they all overlook the inherent bias originating from the foundation models. Specifically, the Web-scale data for pre-training foundation models exhibit a highly skewed distribution due to Zipf’s law of nature reed2001pareto . The resulting foundation models develop a biased decision boundary that leads to a poor zero-shot performance on rare classes. As evidenced in Figure 1(a) and (b), the purple line encounters a dramatic drop, and the zero-shot performance of tail classes is significantly lower than that of head classes ($57.2\%$ vs. $78.0\%$). Existing ensemble methods like WiSE-FT wortsman2022robust overlook the label bias, resulting in an improvement in top-1 ($+0.5\%$) and head accuracy ($+1.7\%$) while a noticeable degradation on the tail performances ($-0.9\%$) in Figure 1(b). Another evidence is that the orange line (WiSE-FT) is below the blue line (fine-tuned models) for rare classes in Figure 1(a).
Figure 1: (a) Per class accuracy of CLIP-ViT/B16 on ImageNet. Class index are sorted using the estimated pre-training label prior. Curves are smoothed for better visualization. (b) Beak-down performance of different models on ImageNet. We equally divide the ImageNet classes into three subgroups, according to the class index. Existing ensemble methods like WiSE-FT wortsman2022robust exhibits a clear performance loss on tail classes, while our GLA stands out for all three subgroups.
We propose Generalized Logit Adjustment (GLA), a simple post-hoc method consisting of two steps: 1) removing the label bias of zero-shot model via estimating the label distribution in the pre-training dataset; 2) ensembling the fine-tuned and debiased zero-shot models. As illustrated in Figure 1 (b), our GLA achieves consistent improvement across all three subgroups, particularly showing a significant gain on tail classes ($+1.5\%$). Despite its simplicity, our GLA has a firm statistical grounding: it is the Bayes optimal classifier given the fine-tuned and zero-shot models, thus consistent for minimizing the error on a class-balanced target distribution (Section 4.2). It is worth noting that removing the bias of foundation models is challenging since the label distribution is often inaccessible due to privacy or copyright concerns. In this work, we only use the downstream labeled data and the zero-shot model to estimate the foundation label bias. Specifically, we formulate the problem by adjusting the margin of the zero-shot models such that the lowest error is achieved on the downstream dataset. This grounding translates into strong empirical performance on real-world datasets, covering few-shot, many-shot, and long-tail learning (Section 5).
The contributions and novelties of this work are summarized as follows:
  • We point out the overlooked label bias in foundation models, which originates from the skewness of pre-training distribution that affects the performance of downstream tasks.
  • We formalize the estimation of the label bias as a constrained optimization problem (Section 3.2) with theoretical justification (Section 4.3). The entire process does not require access to the pre-training dataset, making it practical for fine-tuning scenarios. We present Generalized Logit Adjustment (GLA) method, which ensembles the debiased zero-shot and fine-tuned models, and demonstrate its superiority over conventional fine-tuning and ensembling by proving it’s a Bayes optimal classifier (Section 4.2).
  • We build a comprehensive benchmark for evaluation, which considers three real-world settings and three fine-tuning paradigms. The settings are: 1) many-shot learning with abundant data 2) few-shot learning; and 3) long-tail classification, representing a more challenging scenario that combines many-shot and few-shot data (Section 5). The three fine-tuning paradigms include: 1) end-to-end fine-tuning; 2) linear probing; and 3) prompt tuning (Section 3.1).
  • We demonstrate the efficacy of our proposed method GLA by conducting extensive experiments across various settings and fine-tuning paradigms. We observe 1 to 1.5 pp accuracy gains on ImageNet, large averaged improvement (1.4 to 4.6 pp) on 11 few-shot datasets and 2.4 pp averaged accuracy gains on long-tail datasets. (Section 5).
2 Related Work
Image-text foundation models. Foundation models pre-trained by contrastive objectives have set impressive milestones for image and text representation learning, with CLIP radford2021learning , ALIGN jia2021scaling , CoCa yu2022coca and Flamingo alayrac2022flamingo being the exemplars. Such models exhibit impressive prompt-based zero-shot performance on various image recognition downstream tasks. Our method aims to reduce foundation model biases to boost performance in downstream tasks. While allingham2023simple also addresses word frequency bias, we differ in two key areas: Firstly, we debias zero-shot models using fixed prompts, whereas allingham2023simple refines the prompting process. Secondly, our GLA doesn’t require access to a subset of the pre-training data.
Ensembles. Ensemble methods aim to boost performances by combining multiple networks, which can be either implemented by aggregating model outputs dietterich2000ensemble ; bauer1999empirical ; lakshminarayanan2017simple ; freund1997decision ; pmlr-v180-kumar22a ; yi2023invariant , weight-space ensembling wortsman2022robust ; izmailov2018averaging , or ensemble distillation hinton2015distilling ; NEURIPS2020_18df51b9 . For the adaptation of foundation models, several work propose to ensemble the fine-tuned and zero-shot models for better performance: Wortsman et al. wortsman2022robust ensembles them in weight space; ProGrad and ProReg prograd ; zhu2023debiased propose to fuse them via knowledge distillation. Our GLA is orthogonal to these approaches, as it concentrates on mitigating the biases in foundation models that are detrimental to ensemble models.
Logit adjustment. Logit adjustment menonlong ; tang2020long ; kangdecoupling ; wu2021adversarial ; hong2021disentangling is a post-hoc technique to adjust the biased output of classification networks. Kang et al kangdecoupling proposes an element-wise scaling adjustment for the classifier weight. Tang et al tang2020long removes the projection of features on a global biased direction. Menon menonlong derives the theoretically optimal adjustment from the training distribution. Unlike the those approaches which rely on a transparent training data or class distribution, our GLA can eliminate the class bias without accessing to the pre-training statistics.
3 Methods
3.1 Setup
Task. Consider a classification problem with instances $\mathbf{x}\in\mathcal{X}$ and labels $y\in\mathcal{Y}=[K]=\{1,...,K\}$. We have a zero-shot model $f_{\mathsf{zs}}$ (given below), a downstream dataset $\mathcal{D}_{s}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N_{s}}$ drawn from source distribution $P_{s}$ and a fine-tuned model $f_{\mathsf{ft}}$ (given below) trained on the dataset $\mathcal{D}_{s}$. Give a model $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$ that outputs prediction score, we define the risk of $f$ on the target distribution $P_{t}$ as the mis-classification rate: $\mathcal{R}_{t}(f)=\mathbb{E}_{\mathbf{x},y\sim P_{t}}[y\neq\operatorname*{argmax}_{i}f(\mathbf{x})_{i}]$. Our goal is to learn a model $f_{\mathsf{gla}}$ that best leverages $f_{\mathsf{zs}}$ and $f_{\mathsf{ft}}$ that minimizes the risk $\mathcal{R}_{t}$.
Zero-shot models. We primarily explore CLIP radford2021learning for zero-shot models. CLIP consists of a visual encoder $\Phi_{\mathsf{v}}(\mathbf{x})$ and a text encoder $\Phi_{\mathsf{t}}(\mathbf{t})$, producing $l_{2}$-normalized features from an image $\mathbf{x}$ and a text $\mathbf{t}$ respectively. Zero-shot model $f_{\mathsf{zs}}$ for $K$ classes is enabled by matching image features $\mathbf{v}=\Phi_{\mathsf{v}}(\mathbf{x})$ with classification weights $\mathbf{w}_{k}=\Phi_{\mathsf{t}}(\mathbf{t}_{k})$, where $\mathbf{t}_{k}$ is obtained by extending the class name $\{c_{k}\}$ to a pre-defined prompt, e.g., “a photo of a {$c_{k}$}.”. Additional details are provided in Appendix B.2. The probability of $\mathbf{x}$ being classified as $y$ is defined as:
$P(y|\mathbf{x})=\mathrm{softmax}(f_{\mathsf{zs}}(\mathbf{x}))_{y}=\frac{\exp(\mathbf{v}^{T}\mathbf{w}_{y})}{\sum_{k=1}^{K}\exp(\mathbf{v}^{T}\mathbf{w}_{k})}.$ (1)
Fine-tuned models. Standard fine-tuning initializes the model $f_{\mathsf{ft}}$ with the pre-trained parameters and then solve $f_{\mathsf{ft}}=\operatorname*{argmin}_{f}\mathcal{R}_{s}(f)$ to minimize the risk on downstream dataset. We consider three common variants of fine-tuning: (1) end-to-end, where all parameters of $\Phi_{\mathsf{v}}$ and $\mathbf{w}_{k}$ are updated; (2) linear probing, where only $\mathbf{w}_{k}$ is modified while $\Phi_{\mathsf{v}}$ is fixed; (3) prompt tuning, where the text input $\mathbf{t}_{k}$ is learned, while keeping $\Phi_{\mathsf{v}}$ and $\Phi_{\mathsf{t}}$ freezed. See Appendix B.3 for details on fine-tuning methods.
Notation. Let $P_{p}(y)$, $P_{s}(y)$ and $P_{t}(y)$ be the marginal probability of class $y\in[K]$ for pre-training, source (training) and target (test) distribution, respectively. Let $\pi_{p}$ and $\pi_{s}$ denote the log probabilities of class for pre-training and training distribution, i.e., $\pi_{p}(y)=\log P_{p}(y)$ and $\pi_{s}(y)=\log P_{s}(y)$.
3.2 Generalized Logit Adjustment Framework
Fine-tuned models often yield significant gains compared to zero-shot models, and ensembling them can further improve performance. This leads to a natural question: How should we best leverage the zero-shot and fine-tuned models for the prediction tasks? We attempt to answer by proposing generalized logit adjustment in Definition 1.
Definition 1.
(GLA) The Generalized Logit Adjustment (GLA) model $f_{\mathsf{gla}}$ is defined as follows:
$f_{\mathsf{gla}}(\mathbf{x})=f_{\mathsf{ft}}(\mathbf{x})+f_{\mathsf{zs}}(\mathbf{x})-\pi_{s}-\pi_{p}.$ (2)
In Section 4.2, we prove that given the zero-shot and fine-tuned models, our GLA model is the Bayes optimal classifier and no other combination of the two models can outperform it. Here remains one important question: how could we obtain $\pi_{p}$ as we have no access to the pre-training statistics? We provide the estimation process of $\pi_{p}$ in Eq (4) and postpone the justification in Section 4.3. The entire GLA algorithm consists of two steps which is given as follows:
Figure 2: Illustration of debiasing process on ImageNet validation set. (a) The original distribution of zero-shot outputs; (b) the estimated pre-train distribution $\mathbf{q}$ based on our algorithm; (c) the distribution of debiased zero-shot outputs using estimated $\mathbf{q}$.
Step 1: Estimation of $\pi_{p}$. Let $\mathbf{q}$ be an arbitrary probability simplex over $K$ classes. Given the validation data from $P_{t}$ or the balanced training data, we can estimate the $\hat{\pi}_{p}=\log\mathbf{q}^{*}$ as the constrained optimization problem (proof in Section 4.3):
$\mathbf{q}^{*}=\operatorname*{argmin}_{\mathbf{q}}\mathcal{R}_{t}(f_{\mathsf{zs}}-\log\mathbf{q})$
$\mathrm{s.t.}\ \mathbf{q}_{i}\geq 0,\mathrm{for}\ i\in[K],$
$\quad\sum_{i\in[K]}\mathbf{q}_{i}=1.$ (3)
We constrain the sum of $\mathbf{q}$ to be 1 and ensure that each element is non-negative, guaranteeing that it forms a valid probability distribution. We solve the following Lagrangian problem to find optimal $\mathbf{q}^{*}$:
$\min_{\mathbf{q}}\max_{\lambda_{i}\geq 0,\upsilon}\mathcal{R}_{t}(f_{\mathsf{zs}}-\log\mathbf{q})-\sum_{i}\lambda_{i}\mathbf{q}_{i}+\upsilon(1-\sum_{i\in[K]}\mathbf{q}_{i})$ (4)
Step 2: GLA ensembling. Given the estimated $\hat{\pi}_{p}$ and the known downstream training $\pi_{s}$, we ensemble the zero-shot model $f_{\mathsf{zs}}$ and the fine-tuned model $f_{\mathsf{ft}}$ to get our GLA model $f_{\mathsf{gla}}$ via Eq. (2). We can regard $f_{\mathsf{zs}}-\hat{\pi}_{p}$ and $f_{\mathsf{ft}}-\pi_{s}$ as the debiased zero-shot model (Figure 2(c)) and debiased fine-tuned models, respectively. Our GLA is actually ensembling two debiased models. Note that, different from prograd ; wortsman2022robust , we do not require a hyper-parameter to adjust the contribution of the two models, the optimal solution is to combine them equally (see Section 4.2 for justification).
4 Theoretical Analysis
In this section, we explain why our GLA model best ensembles the zero-shot and fine-tuned models (Section 4.2) and justify the estimation process of the pre-training label distribution $\pi_{p}$ (Section 4.3). We start with some preliminaries on Bayes optimal classifier.
4.1 Preliminaries
Suppose we have a pair $(X,Y)\sim P$ takes values in $\mathcal{X}\times\mathcal{Y}$, where $Y$ is the class label of input $X$.
Definition 2.
The 0-1 error (risk) of a classifier $\hat{y}:\mathcal{X}\rightarrow\mathcal{Y}$ on distribution $P$ is given by:
$\mathcal{R}(\hat{y})=P(Y\neq\hat{y}(X))$ (5)
However, the 0-1 error is non-smooth, one typically minimizes a surrogate loss $\ell$, e.g., cross-entropy: $\ell(f(\mathbf{x}),y)=\log[\sum_{i\in[K]}\exp(f(\mathbf{x})_{i}-f(\mathbf{x})_{y})]$ , where $\hat{y}(\mathbf{x})=\operatorname*{argmax}_{i}f(\mathbf{x})_{i}$. It is known that the cross-entropy loss is Bayes consistent zhang2004statistical , i.e., a nearly optimal minimizer of the cross-entropy loss ($\mathbb{E}_{\mathbf{x},y\sim P}[\ell(f(\mathbf{x}),y)$] is also a nearly optimal optimizer of the mis-classification error ($\mathbb{E}_{\mathbf{x},y\sim P}[y\neq\hat{y}(\mathbf{x})]$).
Definition 3.
The Bayes optimal classifier $y^{*}$ for $P$ given input $\mathbf{x}$ is defined as:
$y^{*}(\mathbf{x})=\operatorname*{argmax}_{y\in\mathcal{Y}}P(y|\mathbf{x})$ (6)
It is called Bayes optimal classifier because on the average no other classifier using the same hypothesis and prior knowledge can outperform it.
Lemma 1.
The Bayes optimal classifier $y^{*}$ for $P$ has lower risk than all classifiers $\hat{y}:\mathcal{X}\rightarrow\mathcal{Y}$.
$\mathcal{R}(y^{*})\leq\mathcal{R}(\hat{y})$ (7)
4.2 Generalized Logit Adjustment Leads to Better Ensembling
Zero-shot and fine-tuned models are complementary. We revisit an empirical phenomena observed in wortsman2022robust Section 5.1: After exploring a series of measures of diversity, covering predictions and features, they find that zero-shot and fine-tuned models have diverse predictions, despite sharing the same backbone. As the two data distribution $P_{s}$ and $P_{p}$ is known to be different, the resulting models leverage different cues to predict: fine-tuned models risk exploiting spurious correlations and in-domain patterns which only hold for downstream dataset taori2020measuring ; arjovsky2020out ; On the other hand, zero-shot CLIP models capture stable correlations across diverse domains and exhibit much higher robustness radford2021learning . For instance, zero-shot models rely on robust features for decisions that can achieve high performance on sketch and adversarial samples, while the fine-tuned models that trained on real images typically fail on these samples, as they rely on spurious correlations that only hold on real images. We formulate the phenomena in the following assumption.
Assumption 1.
Zero-shot and fine-tuned models have diverse predictions:
$(f_{\mathsf{ft}}(\mathbf{x})\perp f_{\mathsf{zs}}(\mathbf{x}))|y.$ (8)
We derive the conditional probability $P_{t}(y|f_{\mathsf{ft}}(\mathbf{x}),f_{\mathsf{zs}}(\mathbf{x}))$ w.r.t. the outputs of $f_{\mathsf{zs}}(\mathbf{x})$ and $f_{\mathsf{ft}}(\mathbf{x})$:
Lemma 2.
For a balanced target distribution [^1] This lemma can be easily extended to imbalanced target distributions (proof in Appendix A). Yet, as most test sets are class-balanced, we focus on the balanced case for brevity., where $P_{t}(y)=1/K$ for all $y\in[K]$, we have:
$P_{t}(y|f_{\mathsf{ft}}(\mathbf{x}),f_{\mathsf{zs}}(\mathbf{x}))=\mathrm{softmax}(\underbrace{f_{\mathsf{ft}}(\mathbf{x})+f_{\mathsf{zs}}(\mathbf{x})-\pi_{s}-\pi_{p}}_{\textstyle{f_{\mathsf{gla}}(\mathbf{x})}})(y)$ (9)
Intuitively, since the zero-shot and fine-tuned models provide diverse predictions, conditioned on the two predictions is equivalent to adding the logits in log space. Additionally, as the target distribution is class-balanced, we need to remove the class bias of two models by subtracting $\pi_{s}$ and $\pi_{p}$. The formal proof is given in Appendix A.1. Note that the RHS of Eq. (9) is exactly the softmax output of our GLA model by Definition 1, which exhibits the following property:
Proposition 1.
Let $g:\mathbb{R}^{K}\times\mathbb{R}^{K}\rightarrow\mathbb{R}^{K}$ be an arbitrary function that ensemble the outputs of $f_{\mathsf{zs}}$ and $f_{\mathsf{ft}}$. Our GLA classifier $f_{\mathsf{gla}}$ has lower risk than any function $f_{g}(\mathbf{x})=g(f_{\mathsf{zs}}(\mathbf{x}),f_{\mathsf{ft}}(\mathbf{x}))$, i.e.
$\mathcal{R}_{t}(f_{\mathsf{gla}})\leq\mathcal{R}_{t}(f_{g}).$ (10)
Proof.
From Lemma 2 and Definition 1, we have:
$\operatorname*{argmax}_{y\in\mathcal{Y}}f_{\mathsf{gla}}(\mathbf{x})_{y}=\operatorname*{argmax}_{y\in\mathcal{Y}}\mathrm{softmax}(f_{\mathsf{ft}}(\mathbf{x})+f_{\mathsf{zs}}(\mathbf{x})-\pi_{s}-\pi_{p})_{y}=\operatorname*{argmax}_{y\in\mathcal{Y}}P_{t}(y|f_{\mathsf{ft}}(\mathbf{x}),f_{\mathsf{zs}}(\mathbf{x})),$ (11)
which means $f_{\mathsf{gla}}$ is the Bayes optimal classifier (see Definition 3) given $f_{\mathsf{ft}}(\mathbf{x})$ and $f_{\mathsf{zs}}(\mathbf{x})$. According to Lemma 1, any other classifier $g(f_{\mathsf{ft}}(\mathbf{x}),f_{\mathsf{zs}}(\mathbf{x}))$ must have higher risk, i.e., $\mathcal{R}_{t}(f_{\mathsf{gla}})\leq\mathcal{R}_{t}(f_{g})$. ∎
Proposition 1 demonstrates that our $f_{\mathsf{gla}}$ model is the best model, as it has the lowest risk on target distribution. Proposition 1 further explains the superiority of $f_{\mathsf{gla}}$ over the fine-tuned model $f_{\mathsf{ft}}$ and the naive ensemble $f_{\mathsf{ens}}(\mathbf{x})=f_{\mathsf{ft}}(\mathbf{x})+f_{\mathsf{zs}}(\mathbf{x})$:
Corollary 1.
$f_{\mathsf{gla}}$ performs better than fine-tuned model $f_{\mathsf{ft}}$ and naive emsembling $f_{\mathsf{ens}}$:
$\mathcal{R}_{t}(f_{\mathsf{gla}})\leq\mathcal{R}_{t}(f_{\mathsf{ft}}),\ \mathcal{R}_{t}(f_{\mathsf{gla}})\leq\mathcal{R}_{t}(f_{\mathsf{ens}})$ (12)
Discussion: when do the GLA models degenerate? Note that there are two equality signs in Eq. (12), indicating that the performance of the GLA model can degenerate to be equivalent to that of the fine-tuned model and naive ensembling in the following two cases.
Case 1: For the first equality, if zero-shot model $f_{\mathsf{zs}}(\mathbf{x})$ provides no further information about $y$ given $f_{\mathsf{ft}}(\mathbf{x})$, i.e., $(y\perp f_{\mathsf{zs}}(\mathbf{x}))|f_{\mathsf{ft}}(\mathbf{x})$, then $P_{t}(y|f_{\mathsf{ft}}(\mathbf{x}),f_{\mathsf{zs}}(\mathbf{x}))$ degenerates to $P_{t}(y|f_{\mathsf{ft}}(\mathbf{x}))$ and the first equality applies. However, in practice, as downstream model and zero-shot model provides diverse predictions, we usually encounter strict inequality, i.e., $\mathcal{R}_{t}(f_{\mathsf{gla}})<\mathcal{R}(f_{\mathsf{ft}})$.
Case 2: The second equality applies when pre-training and downstream training distribution are both class-balanced. In fact, the pre-training dataset for foundation models are known to be highly skewed. Therefore, in most cases, we have $\mathcal{R}_{t}(f_{\mathsf{gla}})<\mathcal{R}_{t}(f_{\mathsf{ens}})$.
In summary, the above two equalities are usually unattainable, which means that theoretically, our GLA model performs better than both the fine-tuned and the naive ensemble models.
4.3 Estimate the label bias of the pre-training dataset
However, $\pi_{p}$ is usually unknown as we have no access to pre-training dataset. In this work, we seek to estimate $\pi_{p}$ using the zero-shot models and the downstream data. Similar to Proposition 1, we have the following proposition says that $f_{\mathsf{zs}}-\pi_{p}$ has lower error on target distribution than any other classifiers that use $f_{\mathsf{zs}}$, see Appendix A.2 for the full proof.
Proposition 2.
Let $h:\mathbb{R}^{K}\rightarrow\mathbb{R}^{K}$ be an arbitrary function that predicts labels using the outputs of the zero-shot model $f_{\mathsf{zs}}(\mathbf{x})$. Let the derived classifier be denoted as $f_{h}(\mathbf{x})=h(f_{\mathsf{zs}}(\mathbf{x}))$. The classifier $f_{\mathsf{zs}}-\pi_{p}$ is better than any $f_{h}(\mathbf{x})$: $\mathcal{R}_{t}(f_{\mathsf{zs}}-\pi_{p})\leq\mathcal{R}_{t}(f_{h}(\mathbf{x}))$.
Let $\mathbf{q}$ be an arbitrary probability simplex over $K$ classes, then we have $\mathcal{R}_{t}(f_{\mathsf{zs}}(\mathbf{x})-\pi_{p})\leq\mathcal{R}_{t}(f_{\mathsf{zs}}(x)-\log\mathbf{q})$. Therefore, we choose to optimize a probability simplex $\mathbf{q}$ over $K$ classes such that the model $f_{\mathsf{zs}}-\log\mathbf{q}$ achieves the minimal empirical risk, as formulated in Eq. (3.2) (the Step 1 of GLA algorithm). Once we obtain the estimated class prior $\hat{\pi}_{p}=\log\mathbf{q}$, we can easily implement the GLA model by ensembling $f_{\mathsf{gla}}(\mathbf{x})=f_{\mathsf{ft}}(\mathbf{x})+f_{\mathsf{zs}}(\mathbf{x})-\pi_{s}-\hat{\pi}_{p}$ (the Step 2 of GLA algorithm).
Figure 3: Estimating label bias of CIFAR-10-LT-IB-10.
Toy experiment. We conducted an experiment to show that the estimated label distribution closely approximates the true one. Specifically, we trained a model with a ResNet32 backbone on the imbalanced CIFAR-10-LT cui2019class dataset with an imbalanced ratio of 10. Subsequently, we used only the test set combined with our proposed method to estimate the label distribution. This procedure simulates scenarios where only downstream data is available and the pre-training data is inaccessible.
Figure 3 reveals a strong alignment between the estimated (orange line) and the actual distributions (blue line), which is further emphasized by a small KL-divergence value of 0.00062. The toy experiment validates the effectiveness of our debiasing method.
Discussion. The $\log\mathbf{q}$ we estimated is not the marginal log-probability of the entire pre-training distribution but the label bias matches the downstream distribution. In the above toy experiment, although training and test sets show different label distributions, their conditional distribution $P(\mathbf{x}|y)$ remains invariant. In this case, our estimate will converge to the actual training label bias. For CLIP models, with diverse pre-training data, some might not align with the downstream domain, potentially compromising the accuracy of the estimation of the entire pre-training distribution.
However, we’d like to point out that removing the label bias of entire pre-training distribution may not optimal for downstream tasks. As a thought experiment, consider a pre-training dataset "sketch" and "photo" styles for "dog" and "cat" samples. Suppose the sample size of "dog” and "cat” is equal but there are more "sketch dogs” than "sketch cats”. This means that even if the overall distribution is balanced, each style isn’t, resulting in biased zero-shot predictions. if we aim to deploy models for the "sketch dogs and cats” domain, adjusting the overall label bias is insufficient. Instead, the optimal label bias should be estimated on the “sketch” distribution. We also provide experiments using LAION-400M dataset in Appendix C.2, illustrating the situation when the downstream data diverges from the pre-training set.
5 Experiments
We evaluate our GLA on three real-world scenarios: many-shot (Section 5.1), few-shot (Section 5.2) and long-tail learning (Section 5.3). We show that our GLA boosts performance on all three settings.
5.1 Many-shot learning
Datasets. We use ImageNet deng2009imagenet and CIFAR100 Krizhevsky09learningmultiple for generic object classification, StanfordCars krause20133d for fine-grained classification, and SUN397 xiao2010sun for scene recognition. See Appendix B.1 for details.
Baselines. We compare GLA against four methods: (1) Zero-shot model, (2) Linear Probing (LP), (3) End-to-End fine-tuning (E2E), and (4) weight ensembling method WiSE-FTwortsman2022robust .
Implementation details. We consider two models: CLIP ViT-B/32 and ViT-B/16. For learning-based models, we fine-tune with AdamW using a cosine annealing learning rate scheduler. We fine-tune for 10 epochs on ImageNet and 20 epochs on other datasets. See Appendix B.3 for further details.
Main results. Table 1 compares our GLA with various baselines. We observe that our GLA can increase the performance of end-to-end fine-tuned models: it achieves $1.5\%$ gains on ImageNet. Compared to WiSE-FT, GLA gains $1.1\%$ top-1 accuracy boost on ImageNet dataset,. Beyond generic object recognition, our method also improves accuracy on the fine-grained dataset (Stanford Cars) and the scene recognition dataset (SUN397), by $0.4\%$ and $0.5\%$, respectively.
Method B/32 B/16
Zero-shot 63.2 68.3
LP 75.8 79.9
E2E 76.3 81.3
WiSE-FT 76.9 81.7
GLA 77.9 82.8
(a) ImageNet
Method B/32 B/16
Zero-shot 64.2 67.2
LP 80.5 83.1
E2E 89.4 91.0
WiSE-FT 90.0 91.3
GLA 90.7 91.9
(b) CIFAR100
Method B/32 B/16
Zero-shot 59.7 64.4
LP 77.1 83.1
E2E 83.7 90.2
WiSE-FT 84.5 90.7
GLA 85.0 91.1
(c) Stanford Cars
Method B/32 B/16
Zero-shot 62.2 64.8
LP 75.6 78.0
E2E 80.0 82.4
WiSE-FT 80.8 82.9
GLA 81.2 83.4
(d) SUN397
Table 1: Accuracy of various methods using CLIP ViT-B/32 and ViT-B/16. LP: linear probe; E2E: end-to-end fine-tuning. Results were obtained using the official implementation from WiSE-FT wortsman2022robust .
Breakdown performance analysis. To analyze the impact of pre-training label bias on fine-tuned and ensemble models, we present the breakdown results on ImageNet using CLIP ViT-B/16, as shown in Table 2. Specifically, we sort the class index using the estimated $\pi_{p}$, and assign the top third of the classes as the head classes, the last third as the tail classes, and the remaining classes as the medium classes. Due to the label bias, the zero-shot tail performance is significantly lower than the head one ($57.2\%$ vs. $78.0\%$). The resulting E2E models are also affected by the bias, with the tail performance being $6.3\%$ lower than the head. Existing ensemble WiSE-FT overlooks the bias, exhibiting noticeable degradation on the tail performances ($-0.9\%$) compared to E2E model, while our GLA stands out for all three subgroups.
Method Head Med. Tail All Zero-shot 78.0 69.8 57.2 68.3 E2E 83.6 83.0 77.3 81.3 WiSE-FT 85.3 83.7 76.4 81.7 GLA 85.2 84.3 78.8 82.8 Table 2: Breakdown results on ImageNet.
Model Source Target ViT-B/32 ViT-B/16 ViT-L/14 $f_{\mathsf{zs}}(\mathbf{x})$ 63.4 68.8 75.6 $f_{\mathsf{zs}}(\mathbf{x})-\hat{\pi}_{p}$ 65.4 69.3 76.3 Table 3: Estimated $\hat{\pi}_{p}$ is transferable across different backbones. $\hat{\pi}_{p}$ is estimated using CLIP ViT-B/32.
Estimated $\pi_{p}$ is transferable across different zero-shot models. The estimated $\pi_{p}$ should be transferable across different zero-shot models if they are trained on the same pre-training dataset. To verify this, we employed a CLIP ViT-B/32 based zero-shot model to estimate $\pi_{p}$, which is subsequently used to debias zero-shot models based on CLIP ViT-B/16 and ViT-L/14. As shown in Table 3, our debiased models outperform the original zero-shot versions by a clear margin.
Figure 4: Accuracy with mixing coefficient $\alpha$.
Ensembling with mixing coefficient. In Section 5.1, we prove the optimal solution is to combine the debiased zero-shot and fine-tuned models equally. We now examine the claim by introducing a mixture coefficient $\alpha\in[0,1]$. The ensemble predictions are given by: $f_{\mathsf{gla}}(\mathbf{x},\alpha)=(1-\alpha)\cdot(f_{\mathsf{zs}}(\mathbf{x})-\pi_{p})+\alpha\cdot(f_{\mathsf{ft}}(\mathbf{x})-\pi_{s})$. We compare the GLA and the naive ensembling with mixture $\alpha$ in Figure 4, where GLA meets its optimal performance at $\alpha=0.5$, which is in line with our theoretical analysis. We also observe that the debiased zero-shot model increases accuracy by $2.3\%$ and our GLA consistently outperforms naive ensembling with various $\alpha$.
5.2 Few-shot learning
For few-shot scenarios, we primarily choose prompt tuning for fine-tuning, since it is empirically more effective than end-to-end fine-tuning and linear probing COOP ; prograd ; PLOT .
Figure 5: Accuracy (%) of few-shot learning on 11 datasets.
Datasets. We follow CoOpCOOP to use 15 datasets: ImageNet deng2009imagenet , Caltech101 fei2004learning , OxfordPets parkhi2012cats , StanfordCars krause20133d , Flowers102 nilsback2008automated , Food101 bossard2014food , FGVCAircraft maji2013fine , EuroSAT helber2019eurosat , UCF101 soomro2012ucf101 , DTD cimpoi2014describing , SUN397 xiao2010sun , ImageNet-{V2 recht2019imagenet , Sketch wang2019learning , A hendrycks2021natural , R hendrycks2021many }. We randomly select {1, 2, 4, 8, 16} shots for training and use the original test set for evaluation. See Appendix B.1 for details.
Method Source Target IN V2 S A R CLIP 59.8 52.8 35.5 22.8 60.6 CoOp 61.9 54.3 32.5 21.8 54.2 PLOT 63.0 55.1 33.0 21.9 55.6 ProGrad 63.5 55.4 33.1 21.3 55.2 GLA 65.6 57.1 36.4 23.0 62.1 Table 4: Evaluation on robustness to distribution shift at 16 training shots.
[Uncaptioned image] Figure 6: Comparison with naive ensembling.
Baselines. We compare with three prompt tuning methods: (1) CoOp COOP optimizes prompts via empirical risk minimization; (2) ProGrad prograd prevents forgetting the general knowledge using zero-shot predictions; (3) PLOT PLOT applies optimal transport to match the vision and text modalities.
Implementation details. We implement our proposed GLA method using CLIP-ResNet-50 as the foundation model and adopt class-specific prompt tuning from CoOp. Results are averaged over three seeds, with training configurations aligned with CoOp. See Appendix B.4 for further information.
Main results. Figure 5 summarizes the few-shot results on 11 datasets. The detailed accuracy and standard deviation are in Appendix C.1. Overall, our GLA clearly outperforms the baselines on average performance by a large margin, e.g., our GLA gains $4.6\%$, $4.8\%$, $3.1\%$, $2.6\%$ and $1.6\%$ performance boost over CoOp at $1,2,4,8,16$ shots. In particular, on ImageNet dataset, we observe a large improvement, e.g., $3.9\%,2.9\%,1.9\%,2.0\%$ and $2.2\%$ over ProGrad at $1,2,4,8,16$ shots. Furthermore, on OxfordPerts, Food101 and SUN397 datasets, our method’s performance remains stable and consistently improves with increasing sample sizes, while the one of the baseline methods fluctuates significantly. In particular, on Food101, baseline models even underperform zero-shot models by a large margin, while our GLA shows clearly better performance than zero-shot models.
Robustness to distribution shifts. Following CoOp, we used the ImageNet at 16 training shots as the source domain and assess the robustness on ImageNet-{V2, Sketch, A, R} datasets. Table 4 summarizes the results, where the prompt tuning baselines perform worse on distribution shifted datasets compared to the zero-shot model, as fine-tuning on limited data misleads the model to learn in-distribution correlations. In comparison, our GLA approach makes the best of both fine-tuned and zero-shot models, thus consistently outperforms other methods on both source and target domains.
GLA improves accuracy over naive ensemble. Figure 6 compares the results between our GLA and naive ensembling. We rank the absolute improvements over fine-tuning baseline at 16 training shots. In summary, our GLA demonstrates superior accuracy gains. It is worth noting that the naive ensembling does not always lead to improvements, e.g., on EuroSAT, SUN, Caltech, UCF and Aircraft, naive ensembling even underperforms fine-tuning baseline.
Model IN. Caltech. Pets. Cars. Flowers. Food. Aircraft. SUN. DTD EuroSAT UCF. Avg.
$f_{\mathsf{zs}}$ 59.8 87.1 85.8 55.6 65.3 77.9 17.1 60.3 42.3 37.5 61.5 59.1
$f_{\mathsf{zs}}-\hat{\pi}_{p}$ 62.3 89.9 85.9 57.6 67.2 78.6 18.4 63.5 43.0 39.0 62.4 60.7
$\Delta$ +2.5 +2.7 +0.1 +2.0 +1.9 +0.7 +1.3 +3.2 +0.7 +1.4 +0.9 +1.6
Table 5: Comparision between zero-shot models ($f_{\mathsf{zs}}$) and debiased zero-shot models ($f_{\mathsf{zs}}-\hat{\pi}_{p}$).
Debiased zero-shot models perform better. We estimate $\pi_{p}$ at 16 shots and compare the original zero-shot models with the debiased zero-shot models in Table 5. It is clear that the debiasing leads to improvement on all 11 datasets, showcasing an average accuracy gain of $1.6\%$.
5.3 Long-tail learning
Datasets and metrics. We evaluate our method on two standard benchmarks: Places365-LT and ImageNet-LT liu2019large . In addition to top-1 accuracy, we report the accuracy on three test subsets according to the number of samples per class: many-shot ($>100$ samples), medium-shot ($20\sim 100$ samples), and few-shot ($<20$ samples). Detailed information is provided in Appendix B.1.
Fine-tuning and long-tail learning baselines. We compare our GLA with the combinations of fine-tuning protocols and long-tailed recognition methods. We consider three fine-tuning protocols: 1) Linear Probe (LP) 2) End-to-End (E2E) fine-tuning; 3) Prompt Tuning (PT). The three fine-tuning paradigms are introduced in Section 3.1 with details in Appendix B.3. We compare with 5 long-tail learning methods: 1) standard Empirical Risk Minimization (ERM); 2) Learnable Weight Scaling (LWS) kangdecoupling ; 3) Logit Adjustment (LA) menonlong ; 4) Balanced Softmax (BS) ren2020balanced , and 5) BALLAD ma2021simple , which is designed for VLMs. See Appendix B.6 for more details on long-tail baselines.
Implementation Details. For all combinations of the fine-tuning and long-tail learning baselines, visual backbones are initialized from CLIP-ResNet-50 and classifiers are initialized via zero-shot prompting. We use SGD for 50 epochs with batch size of 512. See Appendix B.6 for further details.
Results. Table 6 shows that our GLA method consistently surpasses baselines across all long-tailed datasets. Our approach outperforms PT-based models by $3.5\%$ and $4.4\%$ on ImageNet-LT and Places365-LT, respectively. Against E2E approaches, GLA exceeds not just the WiSE-FT but also the current SOTA method BALLAD, by a large margin, e.g., 1pp gains on ImageNet-LT.
Method Many Med Few All
LP ERM 64.4 39.8 20.1 46.6
LA 56.7 50.5 46.4 52.3
BS 57.7 50.5 43.9 52.4
GLA 65.5 60.8 57.7 62.2
PT ERM 66.6 57.7 53.8 60.6
LA 67.9 60.0 58.7 62.9
BS 67.0 62.2 59.4 63.5
GLA 70.7 66.5 64.3 67.8
E2E ERM 75.5 56.0 36.2 60.8
LWS kangdecoupling 70.4 62.6 56.2 64.7
LA menonlong 70.4 62.8 56.9 64.9
BS ren2020balanced 71.2 62.5 56.8 65.1
WiSE-FT wortsman2022robust 70.7 65.0 61.1 66.7
BALLAD ma2021simple 71.0 66.3 59.5 67.2
GLA 72.1 66.4 63.4 68.2
(a) ImageNet-LT
Method Many Med Few All
LP ERM 43.9 21.1 9.0 26.9
LA 39.2 33.7 29.5 34.9
BS 39.6 34.0 28.3 34.9
GLA 43.6 40.2 38.8 41.1
PT ERM 47.7 32.6 24.9 36.5
LA 43.9 40.6 39.4 41.5
BS 43.4 42.5 42.3 42.8
GLA 47.0 47.1 47.8 47.2
E2E ERM 46.3 27.4 12.5 31.3
LWS kangdecoupling 41.2 39.0 32.8 38.6
LA menonlong 42.2 39.3 34.0 39.3
BS ren2020balanced 46.7 44.4 39.5 44.3
WiSE-FT wortsman2022robust 46.0 44.0 43.9 44.7
BALLAD ma2021simple 48.7 44.7 42.2 45.7
GLA 47.3 45.4 45.3 46.1
(b) Places365-LT
Table 6: The performances on ImageNet-LT and Places365-LT.
6 Conclusion and Limitation
In this paper, we identify the label bias in foundation models and underscore its adverse effects on downstream task performance. We propose the Generalized Logit Adjustment (GLA) framework for fine-tuning foundation models, which boosts the performance by effectively eliminating label bias and combining diverse predictions from zero-shot and fine-tuned models. We prove that when presented with zero-shot and fine-tuned models, our GLA is the Bayes optimal classifier for downstream task. Extensive experiments across a diverse range of tasks and fine-tuning framework demonstrate the effectiveness of our approach. We believe that the proposed GLA may partially improve the fairness and credibility of foundation models.
The first limitation is that we only focus on the label bias while other forms of model biases, e.g., representation bias bommasani2021opportunities , cannot be addressed by our algorithm yet. The second limitation is that we primarily focus on enhancing the fine-tuning performance for discriminative models. However, applying our GLA framework to generative models presents challenges. For instance, language generation operates as a Markov process, meaning each output depends on previous ones. This implies it’s not straightforward to estimate the biasedness of a sequence with our GLA, as we only compute the bias in a pre-defined and independent label space.
Acknowledgments
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2021-01-002 and AI Singapore AISG2-RP-2021-022).
Appendix A Proofs
A.1 Proof of Lemma 2
Restated Lemma 0 (Lemma 2).
Let $\pi_{t}$ denote the log probability of the target label distribution $\pi_{t}(y)=\log P_{t}(y)$, we have:
$P_{t}(y|f_{\mathsf{ft}}(\mathbf{x}),f_{\mathsf{zs}}(\mathbf{x}))=\mathrm{softmax}(f_{\mathsf{ft}}(\mathbf{x})+f_{\mathsf{zs}}(\mathbf{x})-\pi_{s}-\pi_{p}+\pi_{t})(y).$ (13)
In class-balanced target distribution, Eq. (13) simplifies to:
$P_{t}(y|f_{\mathsf{ft}}(\mathbf{x}),f_{\mathsf{zs}}(\mathbf{x}))=\mathrm{softmax}(f_{\mathsf{ft}}(\mathbf{x})+f_{\mathsf{zs}}(\mathbf{x})-\pi_{s}-\pi_{p})(y).$ (14)
Proof.
Denote the output $\mathbf{e}=f_{\mathsf{ft}}(\mathbf{x})$ and $\mathbf{z}=f_{\mathsf{zs}}(\mathbf{x})$. We first use the Bayes Rule to decompose $P_{t}(y|\mathbf{e},\mathbf{z})$ into $P_{t}(\mathbf{e},\mathbf{z}|y)$, $P_{t}(y)$ and $P_{t}(\mathbf{e},\mathbf{z})$ in Eq. (15), then rewrite $P_{t}(\mathbf{e},\mathbf{z}|y)$ in Eq. (16) according to Assumption 1. Focusing on label shift problem menonlong ; hong2021disentangling ; lipton2018detecting where $P(\mathbf{x}|y)$ does not change, we derive Eq. (17)
$P_{t}(y|\mathbf{e},\mathbf{z})$ $=\frac{P_{t}(\mathbf{e},\mathbf{z}|y)P_{t}(y)}{P_{t}(\mathbf{e},\mathbf{z})}$ (15)
$=P_{t}(\mathbf{e}|y)P_{t}(\mathbf{z}|y)\frac{P_{t}(y)}{P_{t}(\mathbf{e},\mathbf{z})}$ (16)
$=P_{s}(\mathbf{e}|y)P_{p}(\mathbf{z}|y)\frac{P_{t}(y)}{P_{t}(\mathbf{e},\mathbf{z})}$ (17)
$=\frac{P_{s}(y|\mathbf{e})P_{s}(\mathbf{e})}{P_{s}(y)}\frac{P_{p}(y|\mathbf{z})P_{p}(\mathbf{z})}{P_{p}(y)}\frac{P_{t}(y)}{P_{t}(\mathbf{e},\mathbf{z})}$ (18)
$=\frac{P_{s}(y|\mathbf{e})}{P_{s}(y)}\frac{P_{p}(y|\mathbf{z})}{P_{p}(y)}\frac{P_{s}(\mathbf{e})P_{p}(\mathbf{z})P_{t}(y)}{P_{t}(\mathbf{e},\mathbf{z})}$ (19)
Since $\mathbf{e},\mathbf{z}$ are fixed, we can replace the terms that not rely on $y$ with a constant $C_{1}$ in Eq. (21). We replace $P_{s}(y)=\exp(\log P_{s}(y))=\exp(\pi_{s}(y))$, $P_{p}(y)=\exp(\log P_{p}(y))=\exp(\pi_{p}(y))$ and $P_{t}(y)=\exp(\log P_{t}(y))=\exp(\pi_{t}(y))$. Suppose the underlying class-probabilities $P_{s}(y|\mathbf{e})\propto\exp(\mathbf{e}_{y})$ and $P_{p}(y|\mathbf{z})\propto\exp(\mathbf{z}_{y})$ for $y\in[K]$. Denote the constants $C_{s}$ and $C_{p}$ for normalizing $\exp(\mathbf{e}_{y})$ and $\exp(\mathbf{z}_{y})$ into probabilities, and merge all constants to $C=\frac{C_{1}}{C_{s}C_{p}}$, we get Eq. (23)
$P_{t}(y|\mathbf{e},\mathbf{z})$ $=\frac{P_{s}(y|\mathbf{e})}{P_{s}(y)}\frac{P_{p}(y|\mathbf{z})}{P_{p}(y)}P_{t}(y)C_{1}$ (21)
$=\exp(\mathbf{e}+\mathbf{z}-\pi_{s}-\pi_{p}+\pi_{t})(y)\frac{C_{1}}{C_{s}C_{p}}$ (22)
$=C\cdot\exp(\mathbf{e}+\mathbf{z}-\pi_{s}-\pi_{p}+\pi_{t})(y)$ (23)
Because the summation of $P_{t}(y|\mathbf{e},\mathbf{z})$ is 1, $C=1/\sum_{i\in[K]}\exp(\mathbf{e}+\mathbf{z}-\pi_{s}-\pi_{p}+\pi_{t})(i)$. Therefore, we have:
$P_{t}(y|f_{\mathsf{ft}}(\mathbf{x}),f_{\mathsf{zs}}(\mathbf{x}))$ $=P_{t}(y|\mathbf{e},\mathbf{z})$ (24)
$=\frac{\exp(\mathbf{e}+\mathbf{z}-\pi_{s}-\pi_{p})_{y}}{\sum_{i\in[K]}\exp(\mathbf{e}+\mathbf{z}-\pi_{s}-\pi_{p}+\pi_{t})_{i}}$ (25)
$=\mathrm{softmax}(f_{\mathsf{ft}}(\mathbf{x})+f_{\mathsf{zs}}(\mathbf{x})-\pi_{s}-\pi_{p}+\pi_{t})_{y}$ (26)
In class-balanced target distribution case, $\pi_{t}=\log\frac{1}{K}$ is constant. Since the softmax function is invariant to constant offsets, Eq. (26) simplifies to:
$P_{t}(y|f_{\mathsf{ft}}(\mathbf{x}),f_{\mathsf{zs}}(\mathbf{x}))=\mathrm{softmax}(f_{\mathsf{ft}}(\mathbf{x})+f_{\mathsf{zs}}(\mathbf{x})-\pi_{s}-\pi_{p})_{y}$ (27)
A.2 Proof of Proposition 2
Restated Proposition 0 (Proposition 2).
Suppose that the target distribution $P_{p}$ is class-balanced. Let $h:\mathbb{R}^{K}\rightarrow\mathbb{R}^{K}$ be an arbitrary function that predicts labels using the outputs of the zero-shot model $f_{\mathsf{zs}}(\mathbf{x})$. Let the derived classifier be denoted as $f_{h}(\mathbf{x})=h(f_{\mathsf{zs}}(\mathbf{x}))$. The classifier $f_{\mathsf{zs}}-\pi_{p}$ is better than any $f_{h}(\mathbf{x})$: $\mathcal{R}_{t}(f_{\mathsf{zs}}-\pi_{p})\leq\mathcal{R}_{t}(f_{h}(\mathbf{x}))$.
Proof.
Denote the output $\mathbf{z}=f_{\mathsf{zs}}(\mathbf{x})$. Similar to Eq. (15)-Eq. (26), we have
$P_{t}(y|\mathbf{z})$ $=\frac{P_{t}(\mathbf{z}|y)P_{t}(y)}{P_{t}(\mathbf{z})}$ (28)
$=\frac{P_{p}(\mathbf{z}|y)P_{t}(y)}{P_{t}(\mathbf{z})}$ (29)
$=\frac{P_{p}(y|\mathbf{z})}{P_{p}(y)}\frac{P_{t}(y)}{P_{t}(\mathbf{z})}$ (30)
$=\exp(\mathbf{z}-\pi_{p})(y)/\sum_{i\in[K]}\exp((\mathbf{z}-\pi_{p})(i))$ (31)
$=\mathrm{softmax}(\mathbf{z}-\pi_{p})=\mathrm{softmax}(f_{\mathsf{zs}}(\mathbf{x})-\pi_{p})$ (32)
Therefore, we have:
$\operatorname*{argmax}_{y\in\mathcal{Y}}(f_{\mathsf{zs}}(\mathbf{x})-\pi_{p})_{y}=\operatorname*{argmax}_{y\in\mathcal{Y}}\mathrm{softmax}(f_{\mathsf{zs}}(\mathbf{x})-\pi_{p})_{y}=\operatorname*{argmax}_{y\in\mathcal{Y}}P_{t}(y|f_{\mathsf{zs}}(\mathbf{x}))$ (33)
Again, using Lemma 1, any other classifier $f_{h}(\mathbf{x})$ has higher risk than $f_{\mathsf{zs}}(\mathbf{x})-\pi_{p}$, i.e., $\mathcal{R}_{t}(f_{\mathsf{zs}}-\pi_{p})\leq\mathcal{R}_{t}(f_{h}(\mathbf{x}))$. ∎
Appendix B Experimental Details
B.1 Dataset details
Many-shot and few-shot datasets. For many-shot learning, we use ImageNet, CIFAR100, StanfordCars and SUN397 datasets. For few-shot learning, we evaluate models on 15 datasets. The details of each dataset are presented in Table 7.
Dataset Classes Train size Test size Task
ImageNet 1,000 1.28M 50,000 Object-level
CIFAR100 100 50,000 10,000 Object-level
Caltech101 100 4,128 2,465 Object-level
DTD 47 2,820 1,692 Textures
EuroSAT 10 13,500 8,100 Satellite images
FGVCAircraft 100 3,334 3,333 Fine-grained aircraft
Flowers102 102 4,093 2,463 Fine-grained flowers
Food101 101 50,500 30,300 Fine-grained food
OxfordPets 37 2,944 3,669 Fine-grained pets
StanfordCars 196 6,509 8,041 Fine-grained car
SUN397 397 15,880 19,850 Scene-level
UCF101 101 7,639 3,783 Action
ImageNetV2 1,000 - 10,000 Robustness to collocation
ImageNet-Sketch 1000 - 50,889 Robustness to sketch domain
ImageNet-A 200 - 7,500 Robustness to adversarial attack
ImageNet-R 200 - 30,000 Robustness to multi-domains
Table 7: The detailed statistics of datasets for many-shot and few-shot learning.
Long-tail datasets. We use two standard long-tail benchmarks: Places365-LT and ImageNet-LT liu2019large . The skewness of a long-tailed training set is typically represented by imbalanced ratio, which is defined as $N_{\max}/N_{\min}$. $N_{\max}$ ($N_{\min}$) denotes the largest (smallest) number of instances per class. A larger imbalanced ratio means a more imbalanced training set. The test sets are divided into three splits: many-shot subset contains classes with $>100$ images, medium-shot subset includes classes with $\geq 20$ $\leq 100$ images, and few-shot subset covers classes with $<20$ images. Details are listed in Table 8.
Dataset
Size of
all classes
Size of
many classes
Size of
medium classes
Size of
few classes
Size of
training samples
Imbalanced
ratio
Places365-LT 365 131 163 71 62.5K 996
ImageNet-LT 1000 385 479 136 186K 256
Table 8: Details of long-tailed datasets.
B.2 CLIP zero-shot
We use prompt ensembling of 80 prompts provided by CLIP wortsman2022robust for ImageNet, CIFAR100, and Caltech101 to improve performance, i.e., averaging the text embedding of many captions, e.g.., “a photo of a {$c_{k}$}.” and “an image of a {$c_{k}$}.”. For OxfordPets, StanfordCars, Flowers102, Food101, FGVCAircraft, EuroSAT, UCF101, DTD and SUN397, we use the pre-defined prompt from CoOp COOP .
B.3 Fine-tuned models
End-to-end and linear probe fine-tuning. We follow WiSE-FT wortsman2022robust to implement fine-tuning. We initialize the classifier with the zero-shot classifier and the output of the image encoder $\Phi_{\mathsf{v}}$ is normalized during fine-tuning. We fine-tune for a total of 10 epochs using AdamW loshchilov2018decoupled optimizer with default hyper-parameters $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$ and weight decay $0.1$. We choose a batch size of 512. We use the same data augmentation and cosine-annealing learning rate schedule as wortsman2022robust .
B.4 Prompt tuning.
Prompt tuning like CoOp COOP automates prompt engineering by learning the prompt given few samples from downstream tasks. CoOp provides two options of prompt design: unified prompt that is shared among all classes and class-specific prompt that is different for each class. In this paper, we adopt the class-specific prompt design as the fine-tuned model to implement GLA . In specific, given the word embedding $\mathbf{t}_{k}^{0}$ initialized by zero-shot prompts, we aim to learn a collection of class-specific word embedding $\mathbf{R}=\{\mathbf{r}_{k}\}_{k=1}^{K}$, such that the text input $\mathbf{t}_{k}=\mathbf{t}_{k}^{0}+\mathbf{r}_{k}$ minimizes the empirical risk: $\mathbf{R}^{*}=\operatorname*{argmin}_{\mathbf{R}}\mathbb{E}_{\mathbf{x},y}[y\neq\operatorname*{argmax}_{i}f(x;\mathbf{R})_{i}]$.
We adhere CoOp to use CLIP ResNet-50 as image encoder for few-shot classification. The word embedding $\mathbf{R}$ is initialized from zeros. For the $m$ few-shot classification setting (where $m\in\{1,2,4,8,16\}$), we randomly sample $m$ training and $m$ validation points from the respective full datasets. For all few-shot datasets except ImageNet, the training epoch is set to 200 for 16/8 shots, 100 for 4/2 shots, and 50 for 1 shot. For ImageNet, the epoch is set to 50 for all shots. We fine-tune the prompt with SGD optimizer decayed by the cosine annrealing rule. The base initial learning rate and batch size are set to $10^{-4}$ and $32$. When given an $m$-shot sample setting, we increase the learning rate and batch size by $m$ times simultaneously to accelerate the training speed.
B.5 Estimation of the class prior
To estimate the log-probability of the pre-training distribution $\hat{\pi}_{s}=\log\mathbf{q}$, we utilize the optimization toolkit Cooper gallegoPosada2022cooper from https://github.com/cooper-org/cooper. $\mathbf{q}$ is initialized as a uniformed distribution, $\mathbf{q}(y)=\frac{1}{K}$ for all $y\in[K]$. We use the standard SGD as the primal and dual optimizers for 2000 steps.
B.6 Long-tail learning baselines and training details
We compared with 5 long-tailed classification methods:
  • 1.
    Standard ERM: We learn the model by standard empirical risk minimization on the long-tailed data.
  • 2.
    Learnable Weight Scaling (LWS) kangdecoupling : We first learn the model by standard ERM, then fix the model and learn to re-scale the magnitude of the classifier using class-balanced sampling.
  • 3.
    Logit Adjustment (LA) menonlong : We first learn the model by standard ERM, then compensates the long-tailed distribution by subtracting a class-dependent offset to the model outputs.
  • 4.
    Balanced Softmax (BS) ren2020balanced modifies the Softmax cross-entropy loss which explicitly accommodate the label distribution shift during optimization.
  • 5.
    BALLAD ma2021simple first fine-tunes the vision-language models via contrastive loss on long-tailed data, then freezes the backbone and finally employs an adapter to enhance the representations of tail classes with re-sampling strategies.
For all combinations of the fine-tuning baselines and long-tailed learning methods, visual backbones are initialized from CLIP-ResNet-50 and all classifiers are initialized by feeding prompt with class names to the text encoder. We use SGD for all experiments with a momentum of 0.9 for 50 epochs with batch size of 512. The initial learning rate is set to $1.6\times 10^{-3}$ which is decayed by the cosine annealing rule. To mitigate explosive gradients, we use the warmup learning rate equals to $10^{-5}$ during the first epoch. For the sake of fairness in comparison, all hyper-parameters of baselines are carefully searched using grid search on the validation set.
Appendix C Additional Experiments
C.1 Few-shot learning accuracy
We provide mean and standard deviation in Table 9 in for {1, 2, 4, 8, 16} shots on all 11 few-shot learning datasets.
Dataset 1 shot 2 shots 4 shots 8 shots 16 shots
ImageNet 61.65 $\pm$ 0.15 62.64 $\pm$ 0.01 63.32 $\pm$ 0.07 64.51 $\pm$ 0.09 65.61 $\pm$ 0.03
Caltech101 89.08 $\pm$ 0.09 90.25 $\pm$ 0.25 90.98 $\pm$ 0.43 91.90 $\pm$ 0.21 92.58 $\pm$ 0.42
OxfordPets 87.79 $\pm$ 0.15 87.86 $\pm$ 0.21 88.22 $\pm$ 0.21 88.09 $\pm$ 0.27 89.53 $\pm$ 0.16
StanfordCars 60.00 $\pm$ 0.14 63.10 $\pm$ 0.42 66.25 $\pm$ 0.19 69.87 $\pm$ 0.09 73.95 $\pm$ 0.11
Flowers102 73.45 $\pm$ 0.60 81.00 $\pm$ 0.46 88.31 $\pm$ 0.65 92.89 $\pm$ 0.46 95.41 $\pm$ 0.32
Food101 78.41 $\pm$ 0.07 78.62 $\pm$ 0.07 78.68 $\pm$ 0.06 78.85 $\pm$ 0.19 79.54 $\pm$ 0.47
FGVCAircraft 20.22 $\pm$ 0.59 22.09 $\pm$ 0.37 24.65 $\pm$ 0.85 28.23 $\pm$ 0.44 31.99 $\pm$ 0.50
SUN397 64.29 $\pm$ 0.19 66.32 $\pm$ 0.16 68.01 $\pm$ 0.08 69.99 $\pm$ 0.18 71.64 $\pm$ 0.21
DTD 47.38 $\pm$ 1.23 50.75 $\pm$ 1.46 56.90 $\pm$ 0.20 62.73 $\pm$ 0.80 65.78 $\pm$ 0.49
EuroSAT 56.50 $\pm$ 1.34 67.26 $\pm$ 3.58 72.40 $\pm$ 2.43 77.59 $\pm$ 1.84 84.93 $\pm$ 1.89
UCF101 65.32 $\pm$ 0.17 68.42 $\pm$ 0.81 70.88 $\pm$ 0.50 74.23 $\pm$ 0.24 76.07 $\pm$ 0.03
Table 9: GLA Accuracy (%) with standard deviation of few-shot learning on 11 datasets.
C.2 Experiments on LAION-400M
To support our thought experiment in the discussion of Section 4.3, we use the Open-CLIP ViT-B/16 ilharco_gabriel_2021_5143773 , the first 20k image in LAION-400M datasets and the bias estimation method proposed by allingham2023simple to estimate the expected logits across 8 classes: “dog”, “cat”, “squirrel”, “tiger”, “elephant”, “horse”, “pig” and “bird”. The bias estimation proposed by allingham2023simple provides a good estimation of $\log P(\mathbf{x})$ over the labels under pre-training distribution. Our GLA estimates the label bias matches the downstream domain, we consider two downstream domain styles, i.e., “photo” and “sketch” from DomainNet peng2019moment dataset. For each domain, we randomly sampled 50 images for each class.
Method dog cat squirrel tiger elephant horse pig bird
expected logits by allingham2023simple 0.059 0.039 0.043 0.053 0.043 0.062 0.061 0.028
$\pi_{p}$ by “photo” 0.059 0.041 0.047 0.055 0.048 0.062 0.060 0.033
$\pi_{p}$ by “sketch” 0.059 0.043 0.063 0.061 0.067 0.042 0.047 0.056
Table 10: Comparison among different bias estimation using Open-CLIP-ViT-B/16.
Method Sketch Real Photo
Zero-shot Open-CLIP 92.25 97.00
Debiased by allingham2023simple 89.50 97.50
Debiased by GLA 93.00 97.75
Table 11: Performance on sketch and real photo domain.
We present the expected logits estimated by allingham2023simple , along with the one calculated from “photo” and “sketch” downstream domain data in Table 10. Since the softmax is invariant to constant offsets, i.e., $\mathrm{softmax}(\mathbf{x}+c)=\mathrm{softmax}(\mathbf{x})$, we align the three label bias to yield the same logits on “dog” class by subtracting constant values. We observe that the label bias estimated by the "photo" domain align closely with allingham2023simple due to its close resemblance to the pre-trained domain. Conversely, the "sketch" image style, which significantly differs from the pre-training domain, results in a more pronounced deviation in the pre-trained label bias.
Additionally, we apply the three estimated label biases to debias the zero-shot model and evaluate the classification performance. The results are shown in Table 11, where the superiority of our method becomes evident on “sketch” domain (93.00% vs 92.25%). Applying the label bias from allingham2023simple on the "sketch" domain degrades the model’s performance (from 92.25% to 89.50%). This is attributed to the overall pre-training label bias does not adequately reflect the bias specific to the“sketch” domain.